# Software Performance Engineering for Foundation Model-Powered Software (FMware)

### v1

@@ Abstract
The rise of Foundation Models (FMs) like Large Language Models (LLMs) is revolutionizing software development. Despite the impressive prototypes, transforming FMware into production-ready products demands complex engineering across various domains. A critical but overlooked aspect is performance engineering, which aims at ensuring FMware meets performance goals such as throughput and latency to avoid user dissatisfaction and financial loss. Often, performance considerations are an afterthought, leading to costly optimization efforts post-deployment. FMware’s high computational resource demands highlight the need for efficient hardware use. Continuous performance engineering is essential to prevent degradation. This paper highlights the significance of Software Performance Engineering (SPE) in FMware, identifying four key challenges: cognitive architecture design, communication protocols, tuning and optimization, and deployment. These challenges are based on literature surveys and experiences from developing an in-house FMware system. We discuss problems, current practices, and innovative paths for the software engineering community.
Index Terms:
@@
Foundation Model, Large Language Model, FMware, Software Performance Engineering
I Introduction
@@
The rapid emergence of Foundation Models (FMs), particularly Large Language Models (LLMs), is reshaping software development, with market value expected to reach \$36.1 billion by 2030 [1]. FMs empower the creation of intelligent software, defined as FMware by Hassan et al. [2], where applications rely on one or more building blocks that are FMs.
@@
Many cool demos built with FMware have emerged recently [3, 4]. However, developing FMware from prototypes into production-ready products is a complex engineering process, requiring collaborations across AI, software engineering, systems, and hardware domains throughout the lifetime of such software [5, 6].
@@
Performance engineering, one of the key aspects in such an engineering process, has not been thoroughly discussed. That is, how to proactively ensure that the developed FMware meets the pre-defined performance goals, e.g., throughput or latency. These goals are sometimes also referred to as Service Level Agreements (SLAs) or Service Level Objectives (SLOs). Failing to meet these goals will result in unsatisfactory user experiences.
@@
However, in practice, we observed that performance concerns are often considered afterthoughts during the lifecycle of FMware, causing inefficient and costly performance optimization efforts after the FMware is deployed in production when SLAs are not met. In addition, due to the intensive computation resources that are needed for deploying FMware, it can become prohibitively expensive to serve FMware requests. Efforts to improve the overall efficiency of hardware utilization are needed to avoid the wastage of scarce computing resources, such as costly GPUs sitting idle. Lastly, as FMware is live software that keeps evolving autonomously, it is necessary to apply continuous performance tuning practices to avoid performance degradation over time. To summarize, Software Performance Engineering (SPE) practices are crucial in bringing FMware from prototype to production. Although the awareness of performance-oriented FMware production is growing [5, 6], systematic studies focusing on SPE for FMware (SPE4FMware) are still lacking.
@@
In this paper, we present a comprehensive analysis of SPE challenges in FMware development, deriving from four authoritative sources: (i) an extensive survey of both academic and grey literature, (ii) in-depth discussions with industrial stakeholders and active academicians during SEMLA 2023 & 2024 [7], FM+SE Vision 2030 [8], FM+SE Summit 2024 [9], and SE 2030 workshop - FSE 2024 [10] events, (iii) close collaboration with our customers and our internal FMware application development teams to understand their pain points with performance issues, and (iv) our hands-on experience designing and implementing an in-house FMware serving system (FMware Runtime). We identify four key SPE challenges that span across the lifecycle of FMware development: the design of cognitive architectures, defining communication protocols, tuning and optimization approaches, and deployment options. For each challenge, we describe its aspects in detail, discuss state-of-practices, and share our vision of innovation paths that call for contributions from the software engineering research community.
@@
This paper is organized as follows: Section II outlines the background of our study. Section III delves into the SPE challenges that are associated with FMware. Section IV describes the vision of our serving system. Finally, Section V summarizes our insights and conclusions.
II Background
@@
In this section, we first review SPE research for traditional software (Section II-A). Then we explain the inference process of FM (Section II-B). We also provide an overview of SPE for FM (Section II-C). At last, we present the background of FMware (Section II-D).
II-A Software Performance Engineering (SPE)
@@
SPE involves modeling and analyzing software systems to understand performance characteristics and uncover optimization opportunities [11]. SPE encompasses various engineering practices aimed at meeting performance requirements such as latency, throughput, and resource utilization. As software complexity escalates, addressing performance issues as afterthoughts becomes increasingly challenging and costly [12]. Hence, proactively applying SPE practices and embedding them throughout the development lifecycle is advantageous.
@@
Traditionally, most software components are considered deterministic, allowing developers to recreate issues when diagnosing performance degradation. However, as software systems evolve with increasingly complex interactions, non-deterministic behaviours have emerged [13]. For example, in real-time embedded systems, interrupt-driven interactions with environments introduce unpredictability, as interrupts occur randomly and are handled by priority levels, making overall system behaviour non-deterministic. This non-deterministic behaviour complicates performance engineering, making it difficult to reproduce performance issues that occur in production. This challenge is amplified with FMs due to the probability-based token sampling process in FM inference. Consequently, the rise of FMware necessitates rethinking methods to accurately predict and optimize system performance.
II-B Foundation Models (FMs) & Inference Process
@@
FMs have transformed software by providing unparalleled abilities in comprehending and generating diverse data types. Trained on extensive unlabeled datasets, these models exhibit extraordinary versatility across various tasks, ranging from natural language processing to image generation [14]. Particularly notable are Large Language Models (LLMs), recognized for their sophisticated capabilities in text generation, language comprehension, and multilingual processing. A notable FM architecture is the Generative Pre-trained Transformer (GPT), which employs a decoder-only model architecture, excelling in language understanding and generation. As the number of parameters scales into the billions, the model’s capabilities expand to handle general tasks due to emergent behaviours [15]. A prime example is the renowned GPT-4 model by OpenAI [16].
@@
The inference process of an FM comprises two phases: prefill and decode [17]. In the prefill phase, the user-provided input, or prompt, is fed into the model, initiating the first forward pass to generate the initial output token. This phase is computation-intensive, involving substantial parallel matrix multiplication operations. During the decode phase, models sequentially generate tokens iteratively, where each new token is created based on all previously generated tokens. This token-by-token generation process requires storing the previously computed tokens’ keys and values (known as KV cache) [18] to speed up inference by avoiding redundant computations. The decode phase is memory-bound and cannot be fully parallelized due to data dependency and the sequential nature of token generation.
II-C SPE for FM Inference
@@
The FM inference process exhibits two notable characteristics that impact its performance [19]. First, the queries sent to FM show a diverse length range due to workload heterogeneity. The same task can be articulated through concise instructions or elaborate descriptions, leading to variations in both first token generation latency and KV cache memory consumption. Second, the generated tokens from FM show a diverse length range due to execution unpredictability, resulting in inference completion latency ranging from seconds to minutes and memory consumption varying from megabytes to gigabytes. These characteristics significantly impact how different performance requirements can be satisfied in practice. There are commonly three types of inference tasks: long input and short output (e.g., summarizing an essay), long input and long output (e.g., editing an essay), and short input and long output (e.g., generating an essay). Some tasks mainly demand low latency for the first output token, as subsequent token generation only needs to match human reading speed. In contrast, other tasks require minimal overall latency. These different requirements highlight the importance of performance engineering based on specific use cases.
@@
As FMs continue to expand in size and capability, following the scaling laws [15, 20], optimizing inference becomes essential for efficient FMware deployment. Techniques such as model compression, quantization, and efficient hardware utilization are employed to balance performance with computational demands [21, 22]. A thorough understanding of the FM inference process is critical for advancing SPE for FMware, as FMs serve as the fundamental components. While numerous surveys exist on inference optimization of FMs [21, 23, 22, 24, 25, 26, 27, 28], our paper focuses specifically on the application-level (i.e., FMware) SPE challenges from the perspective of application developers rather than AI engineers – in turn complementing existing efforts for model-level inference optimization.
II-D FM-Powered Software (FMware)
@@
FMs have become pivotal in AI-driven software applications, revolutionizing software engineering and serving as the backbone for a new category of software known as FM-powered software, or FMware [2].
@@
FMware can be classified into two categories: Promptware and Agentware. Promptware involves the direct utilization of FMs through one or many prompts. A notable example is Retrieval-Augmented Generation (RAG)-based software, which enhances output quality by combining FMs with external knowledge sources and documents to avoid issues like hallucination. Promptware varies in complexity, from a single FM invocation in a question-answering session to a meticulously designed chain of invocations represented as Directed Acyclic Graphs (DAG) or workflows [29, 2]. Each node in these workflows represents an individual task, and the edges represent sequential, parallel, or recursive interactions. Tasks can take the form of regular code scripts, traditional ML/DNN model invocations, or FM invocations. This approach enables the creation of compound AI systems capable of handling complex tasks through a series of well-defined steps [30, 2]. Agentware, on the other hand, represents an autonomous and dynamic form of FMware. In Agentware, AI agents powered by FMs can proactively interact with their environment, utilize tools, retain memories, communicate with other AI agents, and autonomously self-explore and improve themselves. While these agents can operate within an explicit workflow similar to Promptware, their true strength lies in autonomy, where researchers expect the agents to reason and develop plans with minimal human intervention. This behaviour emerges during runtime based on interactions and can only be observed through input/output trace data. Although autonomous AI agents remain an active area of research, they are still in the early stages of development and require further exploration [31, 32]. With the characteristics of Promptware and Agentware in mind, we will explore the SPE challenges of FMware in the next section.
III Software Performance Engineering Challenges for FMware
@@
In this section, we describe four challenges in SPE4FMware. For each challenge, we describe the characteristics unique to FMware and introduce a detailed breakdown of the challenge into several dimensions. For each dimension, we present the state of the practices that attempt to tackle the challenge and then discuss the innovation path for future research directions. In particular, the following four challenges are discussed: (1) How to create a high-performance cognitive architecture for FMware (Section III-A)? (2) How to develop a token-efficient communication language among the AI components of an FMware (Section III-B)? (3) How to continuously conduct performance tuning and optimization of FMware (Section III-C)? and (4) How to decide the deployment options for FMware (Section III-D)?
III-A Challenge 1: The complexity of creating high-performance cognitive architectures

### v2

@@ Abstract.
The rise of Foundation Models (FMs) like Large Language Models (LLMs) is revolutionizing software development. Despite the impressive prototypes, transforming FMware into production-ready products demands complex engineering across various domains. A critical but overlooked aspect is performance engineering, which aims at ensuring FMware meets performance goals such as throughput and latency to avoid user dissatisfaction and financial loss. Often, performance considerations are an afterthought, leading to costly optimization efforts post-deployment. FMware’s high computational resource demands highlight the need for efficient hardware use. Continuous performance engineering is essential to prevent degradation. This paper highlights the significance of Software Performance Engineering (SPE) in FMware, identifying four key challenges: cognitive architecture design (i.e., the structural design that defines how AI components interact, reason, and interface with classical software components), communication protocols, tuning and optimization, and deployment. These challenges are based on literature surveys and experiences from developing an in-house FMware system. We discuss problems, current practices, and innovative paths for the software engineering community.
@@
Software Performance Engineering, Foundation Model, FMware, Large Language Model
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software performance
1. Introduction
@@
The rapid emergence of Foundation Models (FMs), particularly Large Language Models (LLMs), is reshaping software development, with market value expected to reach \$36.1 billion by 2030 (Markets and Markets, 2024). FMs empower the creation of intelligent software, which we refer to as FMware (i.e., software built using FMs) (Hassan et al., 2024), where applications rely on one or more building blocks that are FMs.
@@
Demos built with FMware reveal its profound impact across domains such as Software Engineering (e.g., code generation), Knowledge Retrieval (RAG systems), and Autonomous Task Execution, enabling a new class of versatile and intelligent applications (Zhang et al., 2023; Yang et al., 2024). However, developing FMware from prototypes into production-ready products is a complex engineering process, requiring collaborations across AI, systems, software engineering, and hardware domains throughout the lifetime of such software (Yan et al., 2024a; Guo et al., 2024).
@@
Performance engineering, one of the key aspects in such an engineering process, has not been thoroughly discussed. That is, how to proactively ensure that the developed FMware meets the pre-defined performance goals, e.g., throughput or latency. These goals are sometimes also referred to as Service Level Agreements (SLAs) or Service Level Objectives (SLOs). Failing to meet these goals while developing FMware will result in unsatisfactory user experiences or lead to substantial cost.
@@
However, in practice, performance concerns are often considered afterthoughts during the lifecycle of FMware, leading to reactive and costly optimisation efforts once SLAs are not met (Griggs et al., 2024; Fu et al., 2024). Moreover, the intensive computational resources required to serve FMware can make deployment prohibitively expensive. Efforts to improve overall hardware utilization and avoid wastage of scarce computing resources (e.g., idle GPUs) have been highlighted as a critical need in recent work (Fu et al., 2024). Finally, since FMware is live software that evolves continuously in production, continuous performance tuning practices and monitoring are necessary to prevent performance degradation over time (Yan et al., 2024b). To summarize, Software Performance Engineering (SPE) practices are crucial for bringing FMware from prototype to production. Although awareness of performance-oriented FMware production is growing (Yan et al., 2024a; Guo et al., 2024), systematic studies focusing specifically on SPE for FMware (SPE4FMware) remain limited.
@@
In this paper, we present a comprehensive analysis of SPE challenges in FMware development, deriving from various sources by a thorough approach and aiming to reduce any biased opinion. We identify four key SPE challenges that span across the lifecycle of FMware development: the design of cognitive architectures [^1] In this work, we use the term cognitive architecture to refer to the structural design that defines how different AI components interact and reason together to achieve desired outcomes. Establishing a well-structured cognitive architecture is foundational for developing FMware, as it directly affects performance, scalability, and system integration., defining communication protocols, tuning and optimization approaches, and deployment options. For each challenge, we describe its aspects in detail, discuss state-of-practices, and share our vision of innovation paths that call for contributions from the software engineering research community.
@@
This paper is organized as follows: Section 2 outlines the background of our study. Section 3 delves into the SPE challenges that are associated with FMware. Section 4 describes the vision of our serving system. Finally, Section 5 summarizes our insights and conclusions.
2. Background
@@
In this section, we first review Software Performance Engineering (SPE) research for traditional software (Section 2.1). Then we explain the inference process of FM (Section 2.2). We also provide an overview of SPE for FM (Section 2.3). At last, we present the background of FMware (Section 2.4).
2.1. Software Performance Engineering (SPE)
@@
Software Performance Engineering (SPE) was introduced in the early 1980s for systematically addressing performance concerns early and continuously throughout the software development lifecycle (Smith, 1981). Formally, SPE is defined as a quantitative, model-driven approach to predicting, evaluating, and ensuring that software systems meet performance objectives (Smith and Williams, 2002a). Over the decades, SPE has evolved to support object-oriented, distributed, and cloud-native systems, demonstrating its adaptability to emerging software paradigms while maintaining its focus on early prediction and lifecycle-wide performance assurance (Smith and Williams, 2002a). SPE is formalized to involve modeling and analyzing software systems to understand performance characteristics and uncover optimization opportunities (Woodside et al., 2007). SPE encompasses various engineering practices aimed at meeting performance requirements such as latency, throughput, and resource utilization. As software complexity escalates, addressing performance issues as afterthoughts becomes increasingly challenging and costly (Smith and Williams, 2002b). Hence, proactively applying SPE practices and embedding them throughout the development lifecycle is advantageous.
@@
Traditionally, most software components are considered deterministic, allowing developers to recreate issues when diagnosing performance degradation. However, as software systems evolve with increasingly complex interactions, non-deterministic behaviours have emerged (Jain, 1990). For example, in real-time embedded systems, interrupt-driven interactions with environments introduce unpredictability, as interrupts occur randomly and are handled by priority levels, making overall system behaviour non-deterministic. This non-deterministic behaviour complicates performance engineering, making it difficult to reproduce performance issues that occur in production. This challenge is amplified with FMs due to the probability-based token sampling process in FM inference. Consequently, the rise of FMware necessitates rethinking methods to accurately predict and optimize system performance.
2.2. Foundation Models (FMs) & Inference Process
@@
FMs have transformed software by providing unparalleled abilities in comprehending and generating diverse data types. Trained on extensive unlabeled datasets, these models exhibit extraordinary versatility across various tasks, ranging from natural language processing to image generation (Bommasani et al., 2021). Particularly notable are Large Language Models (LLMs), recognized for their sophisticated capabilities in text generation, language comprehension, and multilingual processing. A notable FM architecture is the Generative Pre-trained Transformer (GPT), which employs a decoder-only model architecture, excelling in language understanding and generation. As the number of parameters scales into the billions, the model’s capabilities expand to handle general tasks due to emergent behaviours (Kaplan et al., 2020). A prime example is the renowned GPT-4 model by OpenAI (OpenAI, 2023).
@@
The inference process of an FM comprises two phases: prefill and decode (Agrawal et al., 2023). In the prefill phase, the user-provided input, or prompt, is fed into the model, initiating the first forward pass to generate the initial output token. This phase is computation-intensive, involving substantial parallel matrix multiplication operations. During the decode phase, models sequentially generate tokens iteratively, where each new token is created based on all previously generated tokens. This token-by-token generation process requires storing the previously computed tokens’ keys and values (known as KV cache) (Kwon et al., 2023) to speed up inference by avoiding redundant computations. The decode phase is memory-bound and cannot be fully parallelized due to data dependency and the sequential nature of token generation.
2.3. SPE for FM Inference
@@
The FM inference process exhibits two notable characteristics that impact its performance (Sun et al., 2024a). First, the queries sent to FM show a diverse length range due to workload heterogeneity. The same task can be articulated through concise instructions or elaborate descriptions, leading to variations in both first token generation latency and KV cache memory consumption. Second, the generated tokens from FM show a diverse length range due to execution unpredictability, resulting in inference completion latency ranging from seconds to minutes and memory consumption varying from megabytes to gigabytes. These characteristics significantly impact how different performance requirements can be satisfied in practice. There are commonly three types of inference tasks: long input and short output (e.g., summarizing an essay), long input and long output (e.g., editing an essay), and short input and long output (e.g., generating an essay). Some tasks mainly demand low latency for the first output token, as subsequent token generation only needs to match human reading speed. In contrast, other tasks require minimal overall latency. These different requirements highlight the importance of performance engineering based on specific use cases.
@@
As FMs continue to expand in size and capability, following the scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), optimizing their underlying architecture is essential not only during model training (Narayanan et al., 2021) but also during model inference, as performance optimization is critical for efficient FM deployment. Techniques such as model compression, quantization, and efficient hardware utilization are employed to balance performance with computational demands (Zhou et al., 2024c; Wang et al., 2024c). A thorough understanding of the FM inference process is critical for advancing SPE for FMware, as FMs serve as the fundamental components. While numerous surveys exist on inference optimization of FMs (Zhou et al., 2024c; Yuan et al., 2024a; Wang et al., 2024c; Xu et al., 2024; Liu et al., 2024b; Stojkovic et al., 2024; Li et al., 2024; Bai et al., 2024), our paper focuses specifically on the application-level (i.e., FMware) SPE challenges from the perspective of application developers rather than AI engineers – in turn complementing existing efforts for model-level inference optimization.
2.4. FM-Powered Software (FMware)
@@
FMs have become pivotal in AI-driven software applications, revolutionizing software engineering and serving as the backbone for a new category of software known as FM-powered software, or FMware (Hassan et al., 2024).
@@
FMware can be classified into two categories: Promptware and Agentware. For a visual depiction of this FMware classification, readers may refer to Figure 1 in (Hassan et al., 2024) that illustrates the Promptware and Agentware taxonomy. Promptware refers to a generation of software developed primarily through natural language prompts that directly utilize FMs (Hassan et al., 2024). It involves the direct use of FMs via one or many prompts. A notable example is Retrieval-Augmented Generation (RAG)-based software, which enhances output quality by combining FMs with external knowledge sources and documents to avoid issues like hallucination. Promptware varies in complexity, from a single FM invocation in a question-answering session to a meticulously designed chain of invocations represented as Directed Acyclic Graphs (DAG) or workflows (Wei et al., 2022; Hassan et al., 2024). Each node in these workflows represents an individual task, and the edges represent sequential, parallel, or recursive interactions. Tasks can take the form of regular code scripts, traditional ML/DNN model invocations, or FM invocations. This approach enables the creation of compound AI systems capable of handling complex tasks through a series of well-defined steps (Zaharia et al., 2024; Hassan et al., 2024). Agentware, on the other hand, represents an autonomous and dynamic form of FMware. In Agentware, AI agents powered by FMs can proactively interact with their environment, utilize tools, retain memories, communicate with other AI agents, and autonomously self-explore and improve themselves. While these agents can operate within an explicit workflow similar to Promptware, their true strength lies in autonomy, where researchers expect the agents to reason and develop plans with minimal human intervention. This behaviour emerges during runtime based on interactions and can only be observed through input/output trace data. Although autonomous AI agents remain an active area of research, with recent surveys providing a systematic review of their construction, diverse applications, and evaluation strategies (Wang et al., 2024a), they are still in the early stages of development and require further exploration (Liu et al., 2023; Bouzenia et al., 2024). With the characteristics of Promptware and Agentware in mind, we will explore the SPE challenges of FMware in the next section.
3. Software Performance Engineering Challenges for FMware
@@
In this section, we describe four challenges in SPE4FMware. First, we describe our methodology for deriving such challenges. Then we present the challenges in details. In particular, the following four challenges are discussed: (1) How to create a high-performance cognitive architecture for FMware (Section 3.2)? (2) How to develop a token-efficient communication language among the AI components of an FMware (Section 3.3)? (3) How to continuously conduct performance tuning and optimization of FMware (Section 3.4)? and (4) How to decide the deployment options for FMware (Section 3.5)? Fig. 1 illustrates our four identified SPE challenges and their placement across the FMware development lifecycle. While a comprehensive description of the FMware development lifecycle is detailed in a prior work (Hassan et al., 2024), we utilize this figure to visually anchor each challenge to the specific phases where it primarily manifests. For instance, Challenges 1 and 2 are concentrated in the design, development, and testing phases, while Challenges 3 and 4 are dominant in the testing, deployment, and maintenance phases. For each challenge, we describe the characteristics unique to FMware and introduce a detailed breakdown of the challenge into several dimensions. For each dimension, we present the state of the practices that attempt to tackle the challenge and then discuss the innovation path for future research directions.
Figure 1. The illustration of software performance engineering challenges for FMware throughout the development lifecycle
3.1. Methodology
We apply a systematic approach to derive SPE challenges for FMwaredevelopment. Our methodology for identifying these challenges follows a multifaceted, empirically grounded approach:
  • Literature Review. We conducted a systematic literature review with a search strategy encompassed both peer-reviewed academic literature and carefully selected industry sources to capture both theoretical foundations and practical insights in this rapidly evolving field. We conducted a structured search of academic paper databases (including IEEE Xplore, ACM Digital Library, and Google Scholar) using primary keywords (i.e., “foundation model”, “large language model”) with performance-related concepts (e.g., “performance engineering”, “software performance engineering”, “inference optimization”, “latency”, “cost”, “service level agreement”, “service level objective”, etc.). To capture recent developments in this fast moving field, we supplemented with grey literature from selected arXiv preprints, industry technical reports, white papers, and engineering blogs from leading AI organizations. This mixed approach allowed us to search for foundational academic knowledge with emerging industry practices. The search was limited to the period 2022 to 2024 to ensure relevance in this fast-moving field. We applied a snowballing technique to ensure coverage and excluded papers focused purely on FM-level performance optimization that did not address FMware performance challenges. When citing non-peer-reviewed sources, we carefully qualified their scope and applicability of findings.
  • Community Engagement. We actively participated in and gathered insights from in-depth discussions during premier software engineering venues, including SEMLA 2023 & 2024 (Khomh et al., 2024), FM+SE Vision 2030 (Hassan et al., 2023a), FM+SE Summit 2024 (Hassan et al., 2023b), and SE 2030 workshop - FSE 2024 (Pezzè, 2023) events. At these events, we conducted semi-structured interviews with industrial stakeholders and leading academicians, documenting their perspectives on performance challenges. These semi-structured interviews followed a predefined protocol centered on: 1) the most frequent performance bottlenecks, and 2) the biggest gaps in current performance engineering practices. All insights were systematically documented and reviewed by the authors to identify recurring themes and common pain points across different stakeholder groups.
  • Industrial Communication. We analyzed real-world pain points related to performance through:
    • Structured interviews with our customers (n=10) in the cloud application domain;
    • Regular feedback & debugging sessions with our internal FMware development teams, where we systematically logged all identified performance issues into a shared knowledge base;
    • Systematic documentation of performance-related issues and incidents in a production environment.
  • Industrial Experience. We validated our findings through our hands-on experience in designing and implementing an in-house FMware serving system (FMware Runtime). This provided practical validation of the identified challenges and their implications.
We actively designed our methodology to ensure the generalizability of the identified challenges beyond the specific context of our industrial experience. This was achieved through two critical mechanisms: (1) Data Source Selection: We synthesized insights from four independent streams—literature review, community engagement, industrial communication, and industrial experience—to identify common, foundational principles rather than project-specific performance issues. (2) Focus on Foundational SPE Gaps: The four challenges were ultimately framed as gaps in the SPE practices for the FMware development lifecycle, which are universal issues. While our internal experience provides essential empirical validation, the core challenges were validated and generalized by the common themes identified across the broader academic and industrial sources.
Following data collection, we employed a triangulation and consensus-building process to finalize the challenges. First, themes derived from the literature review were mapped against the pain points identified through community engagement and industrial communication. We then prioritized the identified challenges based on two criteria: (1) Frequency (how often a challenge appeared across all four data sources) and (2) Impact (the severity of the performance degradation as reported, when available). Challenges that were frequent and had high impact formed the basis of the four fundamental SPE challenges presented in this paper, which span the FMware development lifecycle. Our final findings were iteratively refined through reviews and validated among all the authors to ensure comprehensiveness and accuracy.
3.2. Challenge 1: The complexity of creating high-performance cognitive architectures